2. A place-kicker must kick a football from a point 36.0 m from the goal. Half the crowd hopes the ball will clear the crossbar, which is 3.05 m high. When kicked, the ball leaves the ground with a speed of 20.0 m/s at an angle of 53.0° to the horizontal. A. By how much does the ball clear or fall short of clearing the crossbar? B. Does the ball approach the crossbar while still rising or while falling?


1
Expert's answer
2021-10-24T18:22:21-0400

A. From the particle under constant velocity model in the x direction, find the time at which the ball arrives at the goal:

"x_f = x_i + v_{xi}t \\

t = \frac{x_f-x_i}{x_{xi}} \\

= \frac{36.0-0}{20 \times cos(53°)} \\

=2.99 \;s"

From the particle under constant acceleration model in the y direction, find the height of the ball at this time

"y_f =y_i +v_{yi}t + \frac{1}{2}a_yt^2 \\

y_f = 0 + 20 \times sin(53°) \times 2.99 - \frac{1}{2} \times 9.80 \times 2.99 \\

= 3.94 \;m"

Therefore, the ball clears the crossbar by

"3.94 -3.05=0.89 \;m"

Use the particle under constant acceleration model to find the time at which the ball is at its highest point in its trajectory

"v_{vf} =v_{yi} -gt \\

t = \frac{v_{yi}-v_{yf}}{g} \\

= \frac{20.0 \times sin(53.0°) -0}{9.80} \\

= 1.63 \;s"

Because this is earlier than the time at which the ball reaches the goal, the ball clears the goal on its way down.